# GLDiTalker: Speech-Driven 3D Facial Animation with Graph Latent Diffusion Transformer

### v4

@@ 1 Introduction
Speech-driven talking head generation aims to synthesize realistic and synchronized facial motion from speech, with applications in assistive technologies, such as virtual reality and augmented reality Ping et al. (2013); Wohlgenannt et al. (2020). Despite its potential, the task is challenging due to the many-to-many mapping between audio and facial motion, requiring a balance between accurate lip synchronization and natural expression generation.
@@
To achieve realistic speech-driven talking head generation, various methods have been proposed. Traditional approaches Edwards et al. (2016); Taylor et al. (2012); Xu et al. (2013) rely on manual or rule-based designs, which are often costly and inflexible. In contrast, deep learning methods Karras et al. (2017); Peng et al. (2023a) demonstrate stronger adaptability and have made significant progress. VOCA Cudeiro et al. (2019) employs a CNN-based framework to generate cross-identity facial animations, while MeshTalk Richard et al. (2021) leverages U-Net to enhance expression diversity by modeling latent facial expression spaces. Transformer-based models such as FaceFormer Fan et al. (2022) improve temporal and cross-modal alignment through mechanisms like Biased Cross-Modal Multi-Head Attention. Similarly, CodeTalker Xing et al. (2023) introduces a VQ-VAE-based motion prior to reduce the uncertainty in audio-to-motion mapping. Despite these advancements, existing methods share common limitations. Most approaches rely on deterministic models, which generate identical 3D facial motions for the same input conditions. This contradicts real-world behavior, where non-verbal facial cues are inherently nondeterministic Ng et al. (2022). Even when the same individual repeats the same sentence, subtle variations in facial motion naturally occur, reflecting the complexity of human communication. By failing to capture this variability, current methods often suffer from modality inconsistencies, particularly misalignment between audio and mesh. These issues lead to reduced motion diversity and compromised lip-sync accuracy, ultimately undermining the realism and expressiveness of generated animations.
@@ 2.1 Speech-Driven 3D Facial Animation
Speech-driven talking face generation has attracted significant attention and achieved remarkable progress in recent years Karras et al. (2017); Richard et al. (2021); Fan et al. (2022); Xing et al. (2023); Peng et al. (2023a); Wu et al. (2023); Stan et al. (2023); Shen et al. (2025b). Early methods, such as VOCA Cudeiro et al. (2019), introduce CNN-based frameworks for cross-identity face animation, while later works like Mimic Fu et al. (2024) and EmoTalk Peng et al. (2023b) focus on decoupling speech content, style, and emotion. Transformer-based approaches such as FaceFormer Fan et al. (2022) improve audio-visual mapping through attention mechanisms, and CodeTalker Xing et al. (2023) reduces cross-modal uncertainty using discrete motion priors. Recently, FaceDiffuser Stan et al. (2023) applies diffusion mechanisms to enhance the diversity of facial motion generation. Although these methods address various challenges, they still struggle to simultaneously achieve high accuracy and diversity due to limitations in modality alignment and inherent model constraints. To overcome these issues, we propose GLDiTalker, which introduces a quantized space-time diffusion training pipeline to better balance precision and diversity in speech-driven talking face generation.
Figure 2: Overview of the proposed GLDiTalker. In order to enhance the lip-sync accuracy, Stage 1 employs the Spatial Pyramidal SpiralConv Encoder and temporal transformer encoder to sequentially process the input motions to obtain discrete latent facial motions. To enhance the motion diversity, Stage 2 then utilizes a diffusion-based iterative denoising technique to synthesize the latent facial motions from input speech and speaker identity, which is later decoded to a motion mesh sequence by the pretrained frozen decoder of the first stage. $\oplus$” denotes addition and ”N” denotes $N$ identical blocks, each consisting of a SpiralConv layer and a Downsample layer. The heatmaps show the statistical mean of the motion differences between corresponding frames in multiple generation results within the same sequence.
@@ 2.2 Diffusion in Talking Head Animation
Inspired by the diffusion processes in non-equilibrium thermodynamics, Sohl-Dickstein et al. (2015) proposes diffusion, which differs from generative models like GAN Goodfellow et al. (2020), VAE Kingma and Welling (2014), and Flow Kingma and Dhariwal (2018). Diffusion is a Markov process capable of generating high-quality images from Gaussian noise, exhibiting a certain degree of stochasticity Shen and Tang (2024); Shen et al. (2025a). Recently, many 2D talking head generation methods based on diffusion models have achieved highly realistic results Park et al. (2022); Shen et al. (2022); Zhang et al. (2023); Shen et al. (2023); Xu et al. (2024). For example, using audio as a condition, DREAM-Talk Zhang et al. (2023) designs a two-stage diffusion-based framework that generates diverse expressions while maintaining accurate audio-lip synchronization. DiffTalk Shen et al. (2023) further incorporates reference images and facial landmarks, enabling personality-aware synthesis and generalization across identities without fine-tuning. Similarly, VASA-1 Xu et al. (2024) employs diffusion in a latent space to model overall facial dynamics and head movements, effectively capturing subtle facial expressions and head motions. In the context of 3D facial animation, FaceDiffuser Stan et al. (2023) is the first to explicitly apply diffusion for speech-driven 3D facial motion generation. While diffusion methods inherently enhance diversity due to their stochastic nature, they often struggle to achieve high accuracy, particularly in terms of lip synchronization and motion precision. This trade-off between diversity and accuracy remains a significant challenge. To address this, we propose GLDiTalker, which leverages a quantized space-time diffusion training pipeline to balance diversity and accuracy.
@@ 3.1.2 Quantized Space-Time Diffusion Training Pipeline
To solve the problem of failing to simultaneously achieve high lip-sync accuracy and motion diversity caused by modality misalignment, GLDiTalker follows a two-stage pipeline that consists of a Graph Enhanced Quantized Space Learning Stage and a Space-Time Powered Latent Diffusion Stage, as illustrated in Fig. 2. In the first stage, GLDiTalker focuses on quantization to enhance lip-sync accuracy. Specifically, it employs a spatio-temporal VQ-VAE Van Den Oord et al. (2017) based on graph convolution and transformer to model facial motion into a quantized facial motion prior, ensuring precise alignment between audio and facial motion. In the second stage, the focus shifts to diffusion for generating diverse motions. A transformer-based denoising network is adopted for the reverse diffusion process, which iteratively transforms a standard Gaussian distribution into the facial motion prior. This process is conditioned on audio, speaker identity, and the diffusion step, enabling the generation of realistic and expressive facial animations. The animation results can be directly obtained from the denoising outputs by using the decoder from the first stage.
@@ 3.2 Stage 1: Graph Enhanced Quantized Space Learning for Lip-sync Accuracy Improvement
The Spatial Pyramidal SpiralConv Encoder consists of SpiralConv layers, downsampling layers and a fusion layer, as shown in Fig. 3. SpiralConv Gong et al. (2019), a graph-based convolution operator is introduced to process mesh data, which determines the convolution centre and produces a sequence of enumerated centre vertices based on adjacency, followed by the 1-ring vertices, the 2-ring vertices, and so on, until all vertices containing $k$ rings are included.
@@ 3.3 Stage 2: Space-Time Powered Latent Diffusion for Motion Diversity Enhancement
Audio Encoder $E_{a}$ uses the released hubert-base-ls960 version of the HuBERT architecture pre-trained on 960 hours of 16kHz sampled speech audio. The feature extractor, feature projection layer and the initial two layers of the encoder are frozen, while the remaining parameters are set to be trainable, which enables HuBERT to better explore the motion information from audio. Since the facial motions might be captured with a frequency different from that of the speech tokens, we adjust input representation similar to Stan et al. (2023).
@@
where $CA(q,kv)$, $SA(\cdot)$ and $f(\cdot)$ mean Biased Cross-Modal Multi-Head Attention, Biased Causal Multi-Head Self-Attention and feed forward network, respectively; $q$ denotes query input features and $kv$ denotes key and value input features. The alignment mask Fan et al. (2022) in $CA$ makes the motion features only attend to the speech features at the same position.
@@ 4.1 Datasets and Implementations
We conduct abundant experiments on two public 3D facial datasets, BIWI Fanelli et al. (2010) and VOCASET Cudeiro et al. (2019), both of which have 4D face scans along with audio recordings.
BIWI dataset. The BIWI dataset contains 40 sentences spoken by 14 subjects. The recordings are captured at 25 fps with 23370 vertices per mesh. We follow the data splits of the previous work Fan et al. (2022) and only use the emotional data for fair comparisons. Specifically, the training set (BIWI-Train) contains 192 sentences, the validation set (BIWI-Val) contains 24 sentences, and the testing set are divided into two subsets, in which BIWI-Test-A contains 24 sentences spoken by 6 seen subjects during training and BIWI-Test-B contains 32 sentences spoken by 8 unseen subjects during training. BIWI-Test-A is used for both quantitative and qualitative evaluation and BIWI-Test-B is used for qualitative testing.
VOCASET dataset. The VOCASET dataset consists of 480 audio-visual pairs from 12 subjects. The facial motion sequences are captured at 60 fps with about 4 seconds in length. The 3D face meshes in VOCASET are registered to the FLAME topology Li et al. (2017), with 5023 vertices per mesh. Similar to Fan et al. (2022), we adopt the same training (VOCA-Train), validation (VOCA-Val) and testing (VOCA-Test) splits for qualitative testing.
@@ 4.2 Quantitative Evaluation
  • Diversity measures the diversity of the generated facial motions from the same audio input. We calculate this metric across all samples in the BIWI-Test-A following Ren et al. (2023).
@@
Since all subjects in BIWI-Test-A have all been seen in the training dataset, the model can learn the motion styles of the subjects in the testing set, which exhibits a relatively fixed output under speech-driven conditions. In contrast, the subjects in BIWI-Test-B and VOCASET-Test have not been included in the training dataset. Consequently, the model is only able to predict the animation based on the styles of the subjects in the training set. The results may differ significantly from the ground truth, but it may be reasonable to consider without the styles. Therefore, we only quantitatively compare the performance of our GLDiTalker with state-of-the-art methods on BIWI-Test-A, as in previous work Xing et al. (2023) and Stan et al. (2023).
Quantitative evaluations are shown in Tab. 1. Similar to the superior performance of CodeTalker compared to FaceFormer, GLDiTalker outperforms FaceDiffuser in terms of lip movement and upper facial expression motion prediction (0.1383×$\times 10^{-4}$mm and 0.0751×$\times 10^{-5}$mm decrease in LVE and FDD, respectively) by leveraging VQ-VAE to introduce a motion prior that enhances the robustness of the cross-modal mapping from audio to 3D vertex displacements. This strongly proves the effectiveness of our quantized space-time diffusion pipeline. Although TalkingStyle stands out with the lowest LVE and FDD, our approach ranks closely behind, showcasing only a minor difference. Furthermore, the highest Diversity (8.2176×$\times 10^{-4}$mm) demonstrates the superiority of our model in capturing many-to-many mappings between audio and facial motion. The latent diffusion model that starts with random noise and generates latent features through multiple iterations of denoising extremely enhances motion diversity compared to deterministic methods and explicit diffusion model method.
@@ Methods
(×$\times 10^{-4}$mm))
VOCACudeiro et al. (2019) 6.5563 8.1816 0
MeshTalkRichard et al. (2021) 5.9181 5.1025 0
FaceFormerFan et al. (2022) 5.3077 4.6408 0
CodeTalkerXing et al. (2023) 4.7914 4.1170 0
FaceDiffuserStan et al. (2023) 4.7823 3.9225 $5.6421\times 10^{-5}$
TalkingStyleSong et al. (2024) 4.3646 3.8361 0
@@ 4.5.1 Impact of Graph Enhanced Quantized Space Learning Stage
Tab. 3 illustrates that our model achieves obvious superiority in lip-sync accuracy due to the Graph Enhanced Quantized Space Learning Stage with Spatial Pyramidal SpiralConv Encoder. The lowest LVE and FDD (4.6440×$\times 10^{-4}$mm and 3.8474×$\times 10^{-5}$mm in LVE and FDD, respectively) indicate the considerable impact on the overall visual quality of the generated faces and the comprehensibility of lip movements. With regard to the local connectivity of the graph, Spatial Pyramidal SpiralConv Encoder updates the representation of each mesh vertex by fusing the features with those of the neighbouring vertices. This process preserves the topology information of the vertices, thereby encouraging the model to effectively learn the representation of each vertex and the whole graph structure. In contrast, MLP is typically oriented towards regular input data with overly dense connections that do not take into account the structure of the graph.
A series of ablation experiments on kernel size, dilation coefficient, and layer numbers of Spatial Pyramidal SpiralConv Encoder are provided, as shown in Tab. 4. With respect to kernel size, the data in the first and third rows indicates that a large kernel size confers a significant advantage in LVE and FDD (0.2159×$\times 10^{-4}$mm and 0.1852×$\times 10^{-5}$mm decrease in LVE and FDD, respectively). This may be attributed to the fact that larger convolution kernels have larger receptive field, thus enabling the integration of a greater quantity of information. Furthermore, in terms of the dilation coefficients, the third and fifth rows show that larger dilation coefficients lead to better lip synchronisation, although this is associated with a reduction in facial quality performance (0.1305×$\times 10^{-4}$mm decrease in LVE and 0.1074×$\times 10^{-5}$mm increase in FDD). We believe that the effect of lip synchronisation is more significant, so we retain the option of using a large dilation factor. Finally, for the number of layers, two sets of experiments are conducted to demonstrate the importance of the pyramid layer number. With a dilation factor of 1 (no dilation convolution operation), changing the number of layers from 4 to 2 leads to a decline in performance across all metrics (0.1944×$\times 10^{-4}$mm and 0.3487×$\times 10^{-5}$mm increase in LVE and FDD, respectively), as illustrated in the second and third rows; in the case of dilation factor of 2, modifying the number of layers from 4 to 3 also leads to worse performance (0.5208×$\times 10^{-4}$mm and 0.3493×$\times 10^{-5}$mm increase in LVE and FDD, respectively), as shown in the fourth and fifth rows. This suggests that an increase in the layer number can enhance feature extraction capability to comprehensively capture the intrinsic structure and regularity of the sparse vertex data effectively.
@@ 4.5.2 Impact of Space-Time Powered Latent Diffusion Stage
The last column of the Tab. 3 shows that using diffusion in the quantized space of Graph Enhanced Quantized Space Learning Stage leads to significant improvements in terms of motion diversity ($8.2176\times 10^{-4}$mm). With the same input, the deterministic model can only produce a fixed output, which does not reflect the many-to-many correspondence between speech and facial motions in reality. Instead, our approach uses diffusion in the quantized space, which even has an advantage of several orders of magnitude over FaceDiffuser, which adopts diffusion in the explicit space. We also conducted experiments to remove the Noise Encoder $E_{noise}$, which resulted in a performance degradation, as shown in Tab. 5, indicating the need of Noise Encoder.

### v5

@@ 1 Introduction
Speech-driven talking head generation aims to synthesize realistic and synchronized facial motion from speech, with applications in assistive technologies, such as virtual reality and augmented reality ping2013computer; wohlgenannt2020virtual. Despite its potential, the task is challenging due to the many-to-many mapping between audio and facial motion, requiring a balance between accurate lip synchronization and natural expression generation.
@@
To achieve realistic speech-driven talking head generation, various methods have been proposed. Traditional approaches edwards2016jali; taylor2012dynamic; xu2013practical rely on manual or rule-based designs, which are often costly and inflexible. In contrast, deep learning methods karras2017audio; peng2023selftalk demonstrate stronger adaptability and have made significant progress. VOCA cudeiro2019capture employs a CNN-based framework to generate cross-identity facial animations, while MeshTalk richard2021meshtalk leverages U-Net to enhance expression diversity by modeling latent facial expression spaces. Transformer-based models such as FaceFormer fan2022faceformer improve temporal and cross-modal alignment through mechanisms like Biased Cross-Modal Multi-Head Attention. Similarly, CodeTalker xing2023codetalker introduces a VQ-VAE-based motion prior to reduce the uncertainty in audio-to-motion mapping. Despite these advancements, existing methods share common limitations. Most approaches rely on deterministic models, which generate identical 3D facial motions for the same input conditions. This contradicts real-world behavior, where non-verbal facial cues are inherently nondeterministic ng2022learning. Even when the same individual repeats the same sentence, subtle variations in facial motion naturally occur, reflecting the complexity of human communication. By failing to capture this variability, current methods often suffer from modality inconsistencies, particularly misalignment between audio and mesh. These issues lead to reduced motion diversity and compromised lip-sync accuracy, ultimately undermining the realism and expressiveness of generated animations.
@@ 2.1 Speech-Driven 3D Facial Animation
Speech-driven talking face generation has attracted significant attention and achieved remarkable progress in recent years karras2017audio; richard2021meshtalk; fan2022faceformer; xing2023codetalker; peng2023selftalk; wu2023speech; stan2023facediffuser; shen2025long. Early methods, such as VOCA cudeiro2019capture, introduce CNN-based frameworks for cross-identity face animation, while later works like Mimic fu2024mimic and EmoTalk peng2023emotalk focus on decoupling speech content, style, and emotion. Transformer-based approaches such as FaceFormer fan2022faceformer improve audio-visual mapping through attention mechanisms, and CodeTalker xing2023codetalker reduces cross-modal uncertainty using discrete motion priors. Recently, FaceDiffuser stan2023facediffuser applies diffusion mechanisms to enhance the diversity of facial motion generation. Although these methods address various challenges, they still struggle to simultaneously achieve high accuracy and diversity due to limitations in modality alignment and inherent model constraints. To overcome these issues, we propose GLDiTalker, which introduces a quantized space-time diffusion training pipeline to better balance precision and diversity in speech-driven talking face generation.
Figure 2: Overview of the proposed GLDiTalker. In order to enhance the lip-sync accuracy, Stage 1 employs the Spatial Pyramidal SpiralConv Encoder and temporal transformer encoder to sequentially process the input motions to obtain discrete latent facial motions. To enhance the motion diversity, Stage 2 then utilizes a diffusion-based iterative denoising technique to synthesize the latent facial motions from input speech and speaker identity, which is later decoded to a motion mesh sequence by the pretrained frozen decoder of the first stage. $\oplus$” denotes addition and ”N” denotes NN identical blocks, each consisting of a SpiralConv layer and a Downsample layer. The heatmaps show the statistical mean of the motion differences between corresponding frames in multiple generation results within the same sequence.
@@ 2.2 Diffusion in Talking Head Animation
Inspired by the diffusion processes in non-equilibrium thermodynamics, sohl2015deep proposes diffusion, which differs from generative models like GAN goodfellow2020generative, VAE kingma2013auto, and Flow kingma2018glow. Diffusion is a Markov process capable of generating high-quality images from Gaussian noise, exhibiting a certain degree of stochasticity shen2024imagpose; shen2025imagdressing. Recently, many 2D talking head generation methods based on diffusion models have achieved highly realistic results park2022synctalkface; shen2022learning; zhang2023dream; shen2023difftalk; xu2024vasa. For example, using audio as a condition, DREAM-Talk zhang2023dream designs a two-stage diffusion-based framework that generates diverse expressions while maintaining accurate audio-lip synchronization. DiffTalk shen2023difftalk further incorporates reference images and facial landmarks, enabling personality-aware synthesis and generalization across identities without fine-tuning. Similarly, VASA-1 xu2024vasa employs diffusion in a latent space to model overall facial dynamics and head movements, effectively capturing subtle facial expressions and head motions. In the context of 3D facial animation, FaceDiffuser stan2023facediffuser is the first to explicitly apply diffusion for speech-driven 3D facial motion generation. While diffusion methods inherently enhance diversity due to their stochastic nature, they often struggle to achieve high accuracy, particularly in terms of lip synchronization and motion precision. This trade-off between diversity and accuracy remains a significant challenge. To address this, we propose GLDiTalker, which leverages a quantized space-time diffusion training pipeline to balance diversity and accuracy.
@@ 3.1.2 Quantized Space-Time Diffusion Training Pipeline
To solve the problem of failing to simultaneously achieve high lip-sync accuracy and motion diversity caused by modality misalignment, GLDiTalker follows a two-stage pipeline that consists of a Graph Enhanced Quantized Space Learning Stage and a Space-Time Powered Latent Diffusion Stage, as illustrated in Fig. 2. In the first stage, GLDiTalker focuses on quantization to enhance lip-sync accuracy. Specifically, it employs a spatio-temporal VQ-VAE van2017neural based on graph convolution and transformer to model facial motion into a quantized facial motion prior, ensuring precise alignment between audio and facial motion. In the second stage, the focus shifts to diffusion for generating diverse motions. A transformer-based denoising network is adopted for the reverse diffusion process, which iteratively transforms a standard Gaussian distribution into the facial motion prior. This process is conditioned on audio, speaker identity, and the diffusion step, enabling the generation of realistic and expressive facial animations. The animation results can be directly obtained from the denoising outputs by using the decoder from the first stage.
@@ 3.2 Stage 1: Graph Enhanced Quantized Space Learning for Lip-sync Accuracy Improvement
The Spatial Pyramidal SpiralConv Encoder consists of SpiralConv layers, downsampling layers and a fusion layer, as shown in Fig. 3. SpiralConv gong2019spiralnet++, a graph-based convolution operator is introduced to process mesh data, which determines the convolution centre and produces a sequence of enumerated centre vertices based on adjacency, followed by the 1-ring vertices, the 2-ring vertices, and so on, until all vertices containing $k$ rings are included.
@@ 3.3 Stage 2: Space-Time Powered Latent Diffusion for Motion Diversity Enhancement
Audio Encoder $E_{a}$ uses the released hubert-base-ls960 version of the HuBERT architecture pre-trained on 960 hours of 16kHz sampled speech audio. The feature extractor, feature projection layer and the initial two layers of the encoder are frozen, while the remaining parameters are set to be trainable, which enables HuBERT to better explore the motion information from audio. Since the facial motions might be captured with a frequency different from that of the speech tokens, we adjust input representation similar to stan2023facediffuser.
@@
where $CA(q,kv)$, $SA(\cdot)$ and $f(\cdot)$ mean Biased Cross-Modal Multi-Head Attention, Biased Causal Multi-Head Self-Attention and feed forward network, respectively; $q$ denotes query input features and $kv$ denotes key and value input features. The alignment mask fan2022faceformer in $CA$ makes the motion features only attend to the speech features at the same position.
@@ 4.1 Datasets and Implementations
We conduct abundant experiments on two public 3D facial datasets, BIWI fanelli20103 and VOCASET cudeiro2019capture, both of which have 4D face scans along with audio recordings.
BIWI dataset. The BIWI dataset contains 40 sentences spoken by 14 subjects. The recordings are captured at 25 fps with 23370 vertices per mesh. We follow the data splits of the previous work fan2022faceformer and only use the emotional data for fair comparisons. Specifically, the training set (BIWI-Train) contains 192 sentences, the validation set (BIWI-Val) contains 24 sentences, and the testing set are divided into two subsets, in which BIWI-Test-A contains 24 sentences spoken by 6 seen subjects during training and BIWI-Test-B contains 32 sentences spoken by 8 unseen subjects during training. BIWI-Test-A is used for both quantitative and qualitative evaluation and BIWI-Test-B is used for qualitative testing.
VOCASET dataset. The VOCASET dataset consists of 480 audio-visual pairs from 12 subjects. The facial motion sequences are captured at 60 fps with about 4 seconds in length. The 3D face meshes in VOCASET are registered to the FLAME topology FLAME:SiggraphAsia2017, with 5023 vertices per mesh. Similar to fan2022faceformer, we adopt the same training (VOCA-Train), validation (VOCA-Val) and testing (VOCA-Test) splits for qualitative testing.
@@ 4.2 Quantitative Evaluation
  • Diversity measures the diversity of the generated facial motions from the same audio input. We calculate this metric across all samples in the BIWI-Test-A following ren2023diffusion.
@@
Since all subjects in BIWI-Test-A have all been seen in the training dataset, the model can learn the motion styles of the subjects in the testing set, which exhibits a relatively fixed output under speech-driven conditions. In contrast, the subjects in BIWI-Test-B and VOCASET-Test have not been included in the training dataset. Consequently, the model is only able to predict the animation based on the styles of the subjects in the training set. The results may differ significantly from the ground truth, but it may be reasonable to consider without the styles. Therefore, we only quantitatively compare the performance of our GLDiTalker with state-of-the-art methods on BIWI-Test-A, as in previous work xing2023codetalker and stan2023facediffuser.
[…]
@@ Methods
($\times 10^{-4}$$m$))
VOCAcudeiro2019capture 6.5563 8.1816 0
MeshTalkrichard2021meshtalk 5.9181 5.1025 0
FaceFormerfan2022faceformer 5.3077 4.6408 0
CodeTalkerxing2023codetalker 4.7914 4.1170 0
FaceDiffuserstan2023facediffuser 4.7823 3.9225 $5.6421\times 10^{-5}$
TalkingStyletalkingstyle 4.3646 3.8361 0
@@ 4.5.1 Impact of Graph Enhanced Quantized Space Learning Stage
[…]
@@ 4.5.2 Impact of Space-Time Powered Latent Diffusion Stage
The last column of the Tab. 3 shows that using diffusion in the quantized space of Graph Enhanced Quantized Space Learning Stage leads to significant improvements in terms of motion diversity ($8.2176\times 10^{-4}$$m$). With the same input, the deterministic model can only produce a fixed output, which does not reflect the many-to-many correspondence between speech and facial motions in reality. Instead, our approach uses diffusion in the quantized space, which even has an advantage of several orders of magnitude over FaceDiffuser, which adopts diffusion in the explicit space. We also conducted experiments to remove the Noise Encoder $E_{noise}$, which resulted in a performance degradation, as shown in Tab. 5, indicating the need of Noise Encoder.